Design And analysis algorithm course .



Remarks: In all the algorithms, always explain their correctness and analyze their com- plexity. The complexity should be as small as possible. A correct algorithm with large complexity, may not get full credit

Question 2: Give an algorithm that finds the maximum size subarray (the entries may not be contiguous) that forms an increasing sequence.
Answer #1

P = array of length N
M = array of length N + 1

L = 0
for i in range 0 to N-1:
// Binary search for the largest positive j ≤ L
// such that X[M[j]] < X[i]
lo = 1
hi = L
while lo ≤ hi:
mid = ceil((lo+hi)/2)
if X[M[mid]] < X[i]:
lo = mid+1
else:
hi = mid-1

// After searching, lo is 1 greater than the
// length of the longest prefix of X[i]
newL = lo

// The predecessor of X[i] is the last index of
// the subsequence of length newL-1
P[i] = M[newL-1]
M[newL] = i

if newL > L:
// If we found a subsequence longer than any we've
// found yet, update L
L = newL

// Reconstruct the longest increasing subsequence
S = array of length L
k = M[L]
for i in range L-1 to 0:
S[i] = X[k]
k = P[k]

return S